OpenStudy (waheguru):

How do I solve this: Find the equation of a parabola that has x-intercepts -3 and 1 and a y-intercept of 6.

OpenStudy (waheguru):

y=a(x-r)(x-s) I get how to solve but don't get this step 6 = a (0+3)(0-1) Where do the "0"s come from in the equation?

OpenStudy (anonymous):

The x-intercept is the point at which y = 0 (that means the point that lies on the x axis). Likewise, the y-intercept is the point at which x = 0 (the point that lies on the x axis). So the 0 is there because one variable was set to 0.

OpenStudy (waheguru):

So in the equation y = a(x-r)(x-s) the x's are 0 in this case because we are given a y intercept right? But if say y way 6, 1 then in the equation the x's would have both been 1 right?

OpenStudy (anonymous):

You substitute only one variable at a time. So y = a(0-r)(0-s) or 0 = a(x-r)(x-s)

OpenStudy (waheguru):

yea but if the questions said that y was 6,1 then both x's would have been 1 and the y would have been 6 right?

OpenStudy (anonymous):

Yes, that is correct.

OpenStudy (waheguru):

Thanks o_o